OpenStudy (anonymous):

Particle wid unit mass, have the potential V(x) given below. The angular frequency of small oscillations about the minimum of the potential is?

OpenStudy (anonymous):

\[V(x) = a x^{2} + b/x^{2}\]

OpenStudy (anonymous):

options are a) \[\sqrt{8a}\] b) \[\sqrt{8b}\] c) \[\sqrt{8a/b }\] d) \[\sqrt{8b/a}\]

OpenStudy (anonymous):

what is the formula for frequency of a harmonic oscillator, whn potential is given?

OpenStudy (experimentx):

I asked other few guys, they said \( v(x) = ax^2+ \frac b {x2} \) is not harmonic potential

OpenStudy (experimentx):

\[ v(x) = ax^2+ \frac b {x^2} \]

OpenStudy (anonymous):

not exactly harmonic but a simple oscillator, potential is given so my question is, what is angular frequency? @experimentX

OpenStudy (experimentx):

oscillators have frequency ... i think the graph of potential should be like this  for harmonic oscillator

OpenStudy (anonymous):

No, given potential graph is different.... potential cannt be -ve.. u r talking abt different potential

OpenStudy (experimentx):

well ... i guess i can't help you right now then ... i can't remember what principle to apply

OpenStudy (anonymous):

its can be solved by quantum mechanics ..

OpenStudy (experimentx):

Schodinger's equation??

OpenStudy (anonymous):

@kr7210 : first of all , E=-dv/dx find E.. then find F=mE. now write the equation in the form of F=kX(SHM EQUATION)... and find K. and then find omega...

OpenStudy (experimentx):

Hmm .. I think you meant \[ F = -kx = \frac{d}{dx} v(x)\] ?? right??

OpenStudy (anonymous):

oh sorry yes F=-kX..right